I'm not sure if I am doing this correctly. I really appreciate your help!!
Write the given equation in the slope-intercept form and then find the slope and y-intercept of the corresponding line
x - 2y =0
Thanks again!
Val

Answer by stanbon(48516)

The slope-intercept form is y = mx + b
where "m" is the slope and "b" is the intercept.
Solve your equation for "y" as follows:
x-2y=0
2y=x
y=(1/2)x + 0
So, "m" is 1/2 and "b" is 0
Cheers,
Stan H.